OpenStudy (anonymous):

A car slows down from 20m/s to rest in a distance of 85m. What was its acceleration if it is assumed constant?

OpenStudy (anonymous):

this is what I have so far: initial velocity: 20m/s, final velocity: 0m/s, distance = 83m. Since I want to find acceleration I've used this equation: vf^2=vi^2+2ad where (0m/s^2)=(20m/s)^2+2a(85m). My answer was 570m/s^2 but I know this is wrong.. Where did I go wrong?